University of Washington Correspondence Department of Mathematics, Box 354350

INTRADEPARTMENTAL

September 12, 1995

To:  Dave Collingwood, Mathematics Graduate Program Director
From: James A. Morrow, Professor of Mathematics
Subj: Math 598, Seminar in Technology

Math 598, Seminar in Technology, was offered for the first time in Summer Quarter, 1995. Seventeen Students were registered and about five other students or faculty occasionally attended the meetings of the seminar. The students were mostly finishing their second or third years in the graduate program.

Many of the students had some specialized knowledge of computing which they shared with the other students involved. Student participation in the seminar took two forms. A student either gave a talk and presentation on some item of interest to the participants, or constructed a home page for the World Wide Web and put items of interest on the page. Some students did both. The talks were well prepared and well received. The topics discussed were chosen by the students. Most students were interested in details having to do with html, emacs, and LaTeX. Some students spoke about specialized computing with Matlab, Fortran, or C++. Our computer system manager, Don Bovee, gave a talk about his view of the future of computing and answered questions from the audience.

The participants seemed to enjoy the seminar, as attendance remained high throughout the summer. I forgot to ask for written comments on the seminar from the students but I can solicit e-mail responses from them if you think it is worthwhile. I built a homepage

for the seminar (address http://www.math.washington.edu/~morrow/598/598.html). I

am enclosing a hard copy of some of that material. Students in the class were also given a small html primer to get them started on the web.